# Fourier Analysis of Finite Difference Schemes for the Helmholtz Equation: Sharp Estimates and Relative Errors

### v1

@@ Abstract
We propose an approach based on Fourier analysis to wavenumber explicit sharp estimation of absolute and relative errors of finite difference methods for the Helmholtz equation. We use the approach to analyze the classical centred scheme for the Helmholtz equation with a general smooth source term and Dirichlet boundary conditions in 1D. For the Fourier interpolants of the discrete solution with homogeneous (or inhomogeneous) Dirichlet conditions, we show rigorously that the worst case attainable convergence order of the absolute error is $k^{2}h^{2}$ (or $k^{3}h^{2}$) in the $L^{2}$-norm and $k^{3}h^{2}$ (or $k^{4}h^{2}$) in the $H^{1}$-semi-norm, and that of the relative error is $k^{3}h^{2}$ in both $L^{2}$- and $H^{1}$-semi-norms. Even though the classical centred scheme is well-known, it is the first time that such sharp estimates of absolute and relative errors are obtained. We show also that the Fourier analysis approach can be used as a convenient visual tool for evaluating finite difference schemes in presence of source terms, which is beyond the scope of dispersion analysis.
@@ 1 Introduction
The Helmholtz equation $\Delta u+k^{2}u=f$ is a common model of time-harmonic waves in acoustics [20], geophysics[32] and electrical engineering[3], etc. Standard discretization of the Helmholtz equation suffers from the so-called ‘pollution effect’, explained by [10] as “the accuracy of the numerical solution deteriorates with increasing non-dimensional wave number $k$”. It is proved in [22] that the linear finite element method for the 1D problem on $(0,1)$ with $u(0)=0$ and $u^{\prime}(1)-\mathrm{i}ku(1)=0$ has the error in the $H^{1}$-norm bounded from above by $C(hk+h^{2}k^{3})\|f\|_{L^{2}}$ with $C>0$ independent of $h$, $k$ and $f$, where $Ch^{2}k^{3}\|f\|_{L^{2}}$ is called the ‘pollution term’ which indicates the accuracy deterioration with the ‘rule of thumb’ keeping $kh$ constant. Since then, a lot of progress has been made in wavenumber explicit error estimates of finite element methods [29, 40, 45, 11, 27, 34], including the recent extension to heterogeneous media [5, 7, 18, 25] and multiscale methods [28, 6, 31].
@@
While finite element methods are more advanced and powerful, finite difference methods are still used for their simplicity in some applications [35, 19, 30]. But wavenumber explicit error estimates are rarely seen in the literature of finite difference methods [33, 37, 15, 39, 42]. Error estimates of some compact fourth order schemes in 2D are given in [16] based on the assumption that the coefficient matrix of the discretized system is positive definite (which is true only for small wavenumber $k$). In [38], arbitrarily high order and dispersion free schemes in 1D are derived, and an error analysis following the approach of [16] is given. Recently, in [8] it is shown for a dispersion free 3-point scheme in 1D that the maximum-norm error is bounded from above by $(\frac{1}{k}\|f^{\prime\prime}\|_{L^{\infty}}+k^{2}\|f\|_{L^{\infty}})h^{2}$ up to a constant factor.
@@
It has been well noted that the pollution effect is intuitively related to the fact that the numerical solution “has a wavelength that is different from the exact one”, called ‘dispersion[10]. Therefore, [10] advocated “a tool in order to measure quantitatively the dispersion for various of the proposed methods as a measure of the pollution”. “This measure allows to compare the different methods that have been proposed to reduce the pollution and compare their efficiency.” The tool called ‘dispersion analysis[1] had been used before mainly for time-dependent problems [36].
@@
The idea of dispersion analysis is to find Fourier modes e.g. $\mathrm{e}^{\mathrm{i}\boldsymbol{\xi}\cdot\mathbf{x}}$ satisfying the discretized homogeneous Helmholtz equation $\Delta^{h}u^{h}+k^{2}u^{h}=0$ and from the resulting constraint on $\boldsymbol{\xi}$ find the discrete wavenumber $k^{h}:=|\boldsymbol{\xi}|$ as a function of the direction $\mathbf{l}:=\boldsymbol{\xi}/|\boldsymbol{\xi}|$. For example, if the original Helmholtz equation is used instead of the discretized one, then we find the constraint $|\boldsymbol{\xi}|=k$ defining a spherical surface in the space of Fourier frequencies $\boldsymbol{\xi}$. Such Fourier modes $\mathrm{e}^{\mathrm{i}\boldsymbol{\xi}\cdot\mathbf{x}}=\mathrm{e}^{\mathrm{i}k\mathbf{l}\cdot\mathbf{x}}$ are called ‘plane waves’. The ‘phase difference’ $|k-k^{h}|$ as a function of the unit direction $\mathbf{l}$ is a measure of the dispersion error between the original plane wave $\mathrm{e}^{\mathrm{i}k\mathbf{l}\cdot\mathbf{x}}$ and the discretized dispersive ($k^{h}\neq k$) plane wave $\mathrm{e}^{\mathrm{i}k^{h}\mathbf{l}\cdot\mathbf{x}}$.
@@
It is shown in [23] for polynomial finite element methods in 1D that the pollution term is of the same order as the phase difference. That partially explains why reducing the phase difference leads to reducing the pollution effect. Sharp dispersion error estimates are obtained in [2]. By providing a practical criterion, dispersion analysis can also guide optimization of the discretization parameters; see [44] for the continuous interior penalty finite element method and [35, 41, 9, 8] for finite difference methods. Recently, dispersion reduction has been generalized in [26] to unstructured meshes by minimizing the residual of the original plane waves in the discretized equation. A similar approach based on minimizing “average truncation error of plane waves” is adopted in [15] for some high order finite difference methods.
@@
Albeit intuitive, it is not trivial to translate the ‘dispersion error’ into the ‘pollution effect’. As remarked in [21]:“dispersion analysis gives valuable information on several physical phenomena inherent to the discrete solution and thus leads to qualitative insight into the sources of numerical error,” but “it does not yield, by its nature, quantitative statements on the numerical error itself”. That is why so many efforts have been made in wavenumber explicit error estimation, as we mentioned above.
@@
Moreover, dispersion analysis is insufficient because it considers only the zero source problem and the Fourier modes constrained by the dispersion relation, while nonzero source terms and the other Fourier modes stimulated by the source term are not taken into account. We shall see in Section 4 that two dispersion free schemes in 1D can behave differently with source term. Indeed, in [4] a discretization in 1D was proposed with zero truncation error “for as many right hand sides as possible”, not only the zero right hand side. An eigenvalue (Fourier) analysis is carried out in [12] for the Dirichlet boundary value problem with a point source.
@@
Inspired by [12], we develop here an approach to wavenumber explicit error estimation based on Fourier analysis for general smooth source terms. We analyze in detail a classical centred scheme in 1D because it is simple yet not fully understood. But our main purpose is to illustrate the methodology which is applicable in higher dimensions for general linear schemes.
@@
In the same sense, we also give a sharp estimate of relative errors with source term. Wavenumber explicit a priori estimates of relative errors have been much less studied than of absolute errors. Recent progress for finite element methods can be found in [24, 17].
Moreover, by plotting the symbol errors we can see clearly how the error is distributed over Fourier frequencies. In this sense, it becomes a practical tool for comparing different schemes, e.g. those of the same order and those before and after dispersion correction, which complements the dispersion analysis for working with or without source term. This aspect has been briefly demonstrated in [43].
@@ 2.1 Model problem
In case (i) $g_{0}=g_{1}=0$, we assume $f\in H_{0}^{p}(0,1)$ with $p\geq 1$ and $u\in H^{2}(0,1)\cap H_{0}^{1}(0,1)$. In principle, $f\in L^{2}(0,1)$ is sufficient to guarantee that $u\in H^{2}(0,1)\cap H_{0}^{1}(0,1)$. The smoother source $f$ in $H_{0}^{p}$ or even $C_{c}^{\infty}$ can be thought as an arbitrarily close approximation of the general $L^{2}$ source. For the corresponding theory, we refer to [14]. Smoothness of $f$ also facilitates analysis of discretization schemes for (1). For example, [4] used the condition $f\in H^{1}(0,1)$. Here $f\in H_{0}^{p}(0,1)$ and $u\in H^{2}(0,1)\cap H_{0}^{1}(0,1)$ admit odd extension to the domain $(-1,1)$ and the convergent sine series in the corresponding spaces $H^{p}(-1,1)$ and $H^{2}(-1,1)$:
@@ 2.2 Downsampling error
$f^{h}(x_{j})=\sum_{(\xi/\pi)=1}^{N-1}\widehat{f^{h}}(\xi)\sin(\xi x_{j}),\quad\widehat{f^{h}}(\xi):=2\sum_{j=1}^{N-1}h\sin(\xi x_{j})f^{h}(x_{j})=\hat{f}(\xi)+\sum_{s=\pm 1}s\sum_{m=1}^{\infty}\hat{f}(s\xi+2mN\pi),$
@@
Let $u^{h}$ be the finite difference solution defined at the grid points. To calculate the error $e^{h}:=u-u^{h}$ for case (i), we extend $u^{h}$ to the domain $(0,1)$ by the discrete sine transform
@@
$|e_{2}^{h}|_{1}^{2}:=\int_{0}^{1}|(e_{2}^{h})^{\prime}|^{2}~{}\mathrm{d}x=\frac{1}{2}\sum_{(\xi/\pi)=N}^{\infty}\frac{\xi^{2}}{(k^{2}-\xi^{2})^{2}}|\hat{f}(\xi)|^{2}.$
@@
This estimate is only first order in $h$ due to the fact that we only used the low regularity $f_{\mathrm{high}}\in L^{2}(0,1)$. Using higher regularity, $f\in H_{0}^{p}(0,1)$, in a similar way we can derive
@@
Note that the separation $f=f_{\mathrm{low}}+f_{\mathrm{high}}$ depends on the mesh size $h=1/N$. We have $f_{\mathrm{low}}\to f$ and $f_{\mathrm{high}}\to 0$ in $H^{p}(0,1)$ as $N\to\infty$. By Parseval’s identity or orthogonality of the basis, we see that $|f|_{l}^{2}=|f_{\mathrm{low}}|_{l}^{2}+|f_{\mathrm{high}}|_{l}^{2}$ for all $l=0,..,p$.
@@ 3 Analysis of classical 3-point centred scheme
$e^{h}(x)\!=\!A_{0}(x)g_{0}\!+\!A_{1}(x)g_{1},\quad A_{1}(x)\!=\!\csc(k)\sin(kx)\!-\!\csc(k^{h})\sin(k^{h}x),\quad A_{0}(x)=A_{1}(1-x).$ (10)
@@ 3.1 Error for the Dirichlet problem with zero source
$S_{1}:=\frac{1}{2}\frac{\sin k^{h}}{\sin k}+\frac{1}{2}\frac{\sin k}{\sin k^{h}}-\frac{\sin(k^{h}-k)}{k^{h}-k},\quad S_{2}:=-\frac{(\sin k^{h})\cos k}{2k}-\frac{(\sin k)\cos k^{h}}{2k^{h}}+\frac{\sin(k+k^{h})}{k+k^{h}}.$ (11)
Note that
@@
Let $k^{h}:=\frac{2}{h}\arcsin\frac{kh}{2}$ and $c>0$. If $10<\frac{kh}{2}<$ 1, then $\frac{1}{24}k^{3}h^{2}<k^{h}-k<\frac{\pi-2}{8}k^{3}h^{2}$. If in addition $\sigma_{k}:=\min_{\xi\in\pi\mathbb{N}}|k-\xi|>0$ and $k^{3}h^{2}<\frac{8c\sigma_{k}}{\pi-2}$, then $k^{h}-k<c\sigma_{k}$ and $|\sin k^{h}|>\frac{2(1-c)}{\pi}\sigma_{k}$.
@@
Since $x+\frac{x^{3}}{6}<\arcsin x<x+(\frac{\pi}{2}-1)x^{3}$ for $10<x<$ 1, we have $k\!+\!\frac{1}{24}k^{3}h^{2}<k^{h}<k\!+\!\frac{\pi\!-\!2}{8}k^{3}h^{2}<k\!+\!c\sigma_{k}$. It follows that $\min_{\xi\in\pi\mathbb{N}}|\xi\!-\!k^{h}|>\min_{\xi\in\pi\mathbb{N}}|\xi\!-\!k|-|k^{h}\!-\!k|>(1\!-\!c)\sigma_{k}$. So $|\sin k^{h}|\geq\frac{2(1\!-\!c)}{\pi}\sigma_{k}$. ∎
@@
Under the assumptions of Lemma 2 with $10<c<$ 1, for $S_{1}$ in (12) it holds that
@@
Let $k\in((n_{k}^{+}-1)\pi,n_{k}^{+}\pi)$ for an integer $n_{k}^{+}$. By Lemma 2, it holds that $0<k^{h}-k<c\sigma_{k}<\frac{\pi}{2}$. Moreover, $(n_{k}^{+}-1)\pi+\sigma_{k}\leq k<k^{h}<k+c\sigma_{k}\leq n_{k}^{+}\pi-(1-c)\sigma_{k}$. Hence $(n_{k}^{+}-1)\pi<k<k^{h}<n_{k}^{+}\pi$, and $\cot k-\cot k^{h}=(k^{h}-k)\csc^{2}\tilde{k}>k^{h}-k$ for some $\tilde{k}\in(k,k^{h})$. Since
$x-\frac{1}{6}x^{3}<\sin x<x-\frac{1}{8}x^{3},\quad 1-\frac{1}{2}x^{2}<\cos x<1,\quad\text{ for }x\in\left(0,\frac{\pi}{2}\right),$
@@
$S_{1}<\frac{1}{2}(k^{h}-k)^{2}(\csc^{2}k)+1-\left(1-\frac{1}{6}(k^{h}-k)^{2}\right).$
@@
Let $k^{h}:=\frac{2}{h}\arcsin\frac{kh}{2}$ and $S_{2}$ given in (13). If $10<\frac{kh}{2}<$ 1, then it holds that
$-\left(\frac{1}{4k^{2}}+\frac{1}{4k^{3}}\right)(k^{h}-k)^{2}<S_{2}<\left(\frac{1}{4k^{2}}+\frac{1}{4k^{3}}\right)(k^{h}-k)^{2}.$
@@
$\|u\|\geq\left||g_{1}|\cdot\left\|\frac{\sin(kx)}{\sin k}\right\|-|g_{0}|\cdot\left\|\frac{\sin(k(1-x))}{\sin k}\right\|\right|$
@@
Under the assumptions of Lemma 2 with $10<c<$ 1, for $\tilde{S}_{1}$ in (16) it holds that
@@
Let $k^{h}:=\frac{2}{h}\arcsin\frac{kh}{2}$ and $\tilde{S}_{2}$ be given in (17). If $10<\frac{kh}{2}<$ 1, then it holds that
$-\left(\frac{1}{4k^{2}}+\frac{1}{4k^{3}}\right)(k^{h}-k)^{2}<\tilde{S}_{2}<\left(\frac{1}{4k^{2}}+\frac{1}{4k^{3}}\right)(k^{h}-k)^{2}.$

### v2

# Fourier Analysis of Finite Difference Schemes for the Helmholtz Equation in 1D with Dirichlet Conditions: Sharp Estimates and Relative Errorsthanks: 2020 Mathematics Subject Classification: 65N06, 65N12, 65N15

@@ Abstract
We propose an approach based on Fourier analysis to wavenumber explicit sharp estimation of absolute and relative errors of finite difference methods for the Helmholtz equation in 1D with Dirichlet boundary conditions and general source terms. We use the approach to analyze the classical centred scheme. For the Fourier interpolants of the discrete solution with homogeneous (or inhomogeneous) Dirichlet conditions, we show rigorously that the worst case attainable convergence order of the absolute error is $k^{2}h^{2}$ (or $k^{3}h^{2}$) in the $L^{2}$-norm and $k^{3}h^{2}$ (or $k^{4}h^{2}$) in the $H^{1}$-semi-norm, and that of the relative error is $k^{3}h^{2}$ in both $L^{2}$- and $H^{1}$-semi-norms. Even though the classical centred scheme is well-known, it is the first time that such sharp estimates of absolute and relative errors are obtained. We show also that the Fourier analysis approach can be used as a convenient visual tool for evaluating finite difference schemes in presence of source terms, which is beyond the scope of dispersion analysis.
@@ 1 Introduction
The Helmholtz equation $\Delta u+k^{2}u=f$ is a common model of time-harmonic waves in acoustics [24], geophysics[37] and electrical engineering[4], etc. Standard discretization of the Helmholtz equation suffers from the so-called “pollution effect”, explained by [11] as “the accuracy of the numerical solution deteriorates with increasing non-dimensional wave number kk”. It is proved in [26] that the linear finite element method for the 1D problem on $(0,1)$ with $u(0)=0$ and $u^{\prime}(1)-\mathrm{i}ku(1)=0$ has the error in the $H^{1}$-norm bounded from above by $C(hk+h^{2}k^{3})\|f\|_{L^{2}}$ with $C>0$ independent of $h$, $k$ and $f$, where $Ch^{2}k^{3}\|f\|_{L^{2}}$ is called the “pollution term” which indicates the accuracy deterioration with the “rule of thumb” keeping $kh$ constant. Since then, a lot of progress has been made in wavenumber explicit error estimates of finite element methods [34, 46, 51, 12, 32, 39], including the recent extension to heterogeneous media [6, 8, 21, 30] and multiscale methods [36, 18, 22, 7, 33].
@@
While finite element methods are more advanced and powerful, finite difference methods are still used for their simplicity in some applications [40, 23, 2, 41, 35]. In particular, it was found in [41] for full waveform inversion that “a 27-point stencil on a regular Cartesian grid” is “more computationally efficient” than “a $P_{3}$ finite element method on $h$-adaptive tetrahedral mesh”. It is pressing to have a theoretical support for methods in use. But wavenumber explicit error estimates are rarely seen in the literature of finite difference methods [38, 43, 16, 45, 48]. Error estimates of some compact fourth order schemes in 2D are given in [19] based on the assumption that the coefficient matrix of the discretized system is positive definite (which is true only for small wavenumber $k$). In [44], arbitrarily high order and dispersion free schemes in 1D are derived, and an error analysis following the approach of [19] is given. Recently, in [9] it is shown for a dispersion free 3-point scheme in 1D that the maximum-norm error is bounded from above by $(\frac{1}{k}\|f^{\prime\prime}\|_{L^{\infty}}+k^{2}\|f\|_{L^{\infty}})h^{2}$ up to a constant factor. An earlier dispersion free scheme in 1D was proposed in the seminal paper [5] and analyzed in the finite element framework giving the upper bound of the $H^{1}$-semi-norm error as $ch\|f\|_{H^{1}}$ for the scattering problem (transparent on one end and Dirichlet on the other) with $c$ seemingly $O(k)$.
@@
It has been well noted that the pollution effect is intuitively related to the fact that the numerical solution “has a wavelength that is different from the exact one”, called “dispersion[11]. Therefore, [11] advocated “a tool in order to measure quantitatively the dispersion for various of the proposed methods as a measure of the pollution”. “This measure allows to compare the different methods that have been proposed to reduce the pollution and compare their efficiency.” The tool called “dispersion analysis[1] had been used before mainly for time-dependent problems [42].
@@
The idea of dispersion analysis is to find Fourier modes e.g. $\mathrm{e}^{\mathrm{i}\boldsymbol{\xi}\cdot\mathbf{x}}$ satisfying the discretized homogeneous Helmholtz equation $\Delta^{h}u^{h}+k^{2}u^{h}=0$ and from the resulting constraint on $\boldsymbol{\xi}$ find the discrete wavenumber $k^{h}:=|\boldsymbol{\xi}|$ as a function of the direction $\mathbf{l}:=\boldsymbol{\xi}/|\boldsymbol{\xi}|$. For example, if the original Helmholtz equation is used instead of the discretized one, then we find the constraint $|\boldsymbol{\xi}|=k$ defining a spherical surface in the space of Fourier frequencies $\boldsymbol{\xi}$. Such Fourier modes $\mathrm{e}^{\mathrm{i}\boldsymbol{\xi}\cdot\mathbf{x}}=\mathrm{e}^{\mathrm{i}k\mathbf{l}\cdot\mathbf{x}}$ are called “plane waves”. The “phase difference” $|k-k^{h}|$ as a function of the unit direction $\mathbf{l}$ is a measure of the dispersion error between the original plane wave $\mathrm{e}^{\mathrm{i}k\mathbf{l}\cdot\mathbf{x}}$ and the discretized dispersive ($k^{h}\neq k$) plane wave $\mathrm{e}^{\mathrm{i}k^{h}\mathbf{l}\cdot\mathbf{x}}$.
@@
It is shown in [27] for polynomial finite element methods in 1D that the pollution term is of the same order as the phase difference. That partially explains why reducing the phase difference leads to reducing the pollution effect. Sharp dispersion error estimates are obtained in [3]. By providing a practical criterion, dispersion analysis can also guide optimization of the discretization parameters; see [50] for the continuous interior penalty finite element method and [40, 47, 10, 9] for finite difference methods. Recently, dispersion reduction has been generalized in [31] to unstructured meshes by minimizing the residual of the original plane waves in the discretized equation; see also the earlier work [17]. A similar approach based on minimizing “average truncation error of plane waves” is adopted in [16] for some high order finite difference methods.
@@
Albeit intuitive, it is not trivial to translate the “dispersion error” into the “pollution effect”. As remarked in [25]:“dispersion analysis gives valuable information on several physical phenomena inherent to the discrete solution and thus leads to qualitative insight into the sources of numerical error,” but “it does not yield, by its nature, quantitative statements on the numerical error itself”. That is why so many efforts have been made in wavenumber explicit error estimation, as we mentioned above.
@@
Moreover, dispersion analysis is insufficient because it considers only the zero source problem and the Fourier modes constrained by the dispersion relation, while nonzero source terms and the other Fourier modes stimulated by the source term are not taken into account. We shall see in Section 4 that three dispersion free schemes in 1D can behave differently with source term. Indeed, in [5] a discretization in 1D was proposed with zero truncation error “for as many right hand sides as possible”, not only the zero right hand side. An eigenvalue (Fourier) analysis is carried out in [13] for the Dirichlet boundary value problem with a point source.
@@
Inspired by [13] which calculated the eigenvalue difference of the continuous and discrete Helmholtz operators, we develop here an approach to wavenumber explicit error estimation based on Fourier analysis for general smooth source terms and Dirichlet boundary conditions in 1D. We analyze in detail a classical centred scheme in 1D because it is simple yet not fully understood. Actually, we have not found any other wavenumber explicit estimate for finite difference schemes than that in [9], and in particular none for the classical centred scheme, though the folklore says a $p$th order scheme has the (relative) error $O(k^{p+1}h^{p})$ (see e.g. [13]). So we believe our main results (see the Abstract for a preview) are novel.
@@
We also give a sharp estimate of relative errors with source term. Wavenumber explicit a priori estimates of relative errors have been much less studied than of absolute errors. Recent progress for finite element methods can be found in [29, 20].
Moreover, by plotting the symbol errors we can see clearly how the error is distributed over Fourier frequencies. In this sense, it becomes a practical tool for comparing different schemes, e.g. those of the same order and those before and after dispersion correction, which complements the dispersion analysis for working with or without source term. This aspect has been briefly demonstrated in [49] for some compact high order schemes, where sharp orders in $k$, $h$ are found in Fourier domain and verified numerically.
@@ 2.1 Model problem
In case (i) $g_{0}=g_{1}=0$, we assume $f\in H_{0}^{p}(0,1)$ with $p\geq 1$ and $u\in H^{2}(0,1)\cap H_{0}^{1}(0,1)$. In principle, $f\in L^{2}(0,1)$ is sufficient to guarantee that $u\in H^{2}(0,1)\cap H_{0}^{1}(0,1)$. The smoother source $f$ in $H_{0}^{p}$ or even $C_{c}^{\infty}$ can be thought as an arbitrarily close approximation of the general $L^{2}$ source. For the corresponding theory, we refer to [15]. Smoothness of $f$ also facilitates analysis of discretization schemes for (1). For example, [5] used the condition $f\in H^{1}(0,1)$. Here $f\in H_{0}^{p}(0,1)$ and $u\in H^{2}(0,1)\cap H_{0}^{1}(0,1)$ admit odd extension to the domain (1,1)(-1,1) and the convergent sine series in the corresponding spaces $H^{p}(-1,1)$ and $H^{2}(-1,1)$:
@@ 2.2 Downsampling error
$f^{h}(x_{j})=\sum_{(\xi/\pi)=1}^{N-1}\widehat{f^{h}}(\xi)\sin(\xi x_{j}),\quad\widehat{f^{h}}(\xi)=2\sum_{j=1}^{N-1}h\sin(\xi x_{j})f^{h}(x_{j})=\hat{f}(\xi)+\sum_{s=\pm 1}s\sum_{m=1}^{\infty}\hat{f}(s\xi+2mN\pi),$
@@
Let $u^{h}$ be the finite difference solution defined at the grid points. To calculate the error $e^{h}:=u-u^{h}$ for case (i), we extend $u^{h}$ to the domain $(0,1)$ by the discrete sine transform [^1] The sine interpolation is chosen to facilitate our analysis because the continuous solution $u$ is expanded to sine series. In [5], piecewise linear interpolation is used for analysis of a linear finite element method.
@@
$|e_{2}^{h}|_{1}^{2}:=\int_{0}^{1}|(e_{2}^{h})^{\prime}|^{2}~\mathrm{d}x=\frac{1}{2}\sum_{(\xi/\pi)=N}^{\infty}\frac{\xi^{2}}{(k^{2}-\xi^{2})^{2}}|\hat{f}(\xi)|^{2}.$ (5)
@@
Using higher regularity $f\in H_{0}^{p}(0,1)$ allows the split $\frac{\xi^{2}}{(k^{2}-\xi^{2})^{2}}\frac{1}{\xi^{2p}}|\xi^{p}\hat{f}(\xi)|^{2}$ in (5) and yields
@@
Note that the separation $f=f_{\mathrm{low}}+f_{\mathrm{high}}$ depends on the mesh size $h=1/N$. We have $f_{\mathrm{low}}\to f$ and $f_{\mathrm{high}}\to 0$ in $H^{p}(0,1)$ as $N\to\infty$. By Parseval’s identity or orthogonality of the basis, we see that $|f|_{l}^{2}=|f_{\mathrm{low}}|_{l}^{2}+|f_{\mathrm{high}}|_{l}^{2}$ for all $l=0,..,p$. If $|f_{\mathrm{high}}|_{p}=O(k^{p})$ and $kh=O(1)$, then (7) and (8) say $|e_{2}^{h}|_{1}=O(h)$ and $\|e_{2}^{h}\|=O(h^{2})$, so the higher regularity does not improve the estimates.
@@ 3 Analysis of classical 3-point centred scheme
$e^{h}(x)\!=\!A_{0}(x)g_{0}\!+\!A_{1}(x)g_{1},\quad A_{1}(x)\!=\!\frac{\sin(kx)}{\sin k}\!-\!\frac{\sin(k^{h}x)}{\sin k^{h}},\quad A_{0}(x)=A_{1}(1-x).$ (11)
@@ 3.1 Error for the Dirichlet problem with zero source
$S_{1}:=\frac{1}{2}\frac{\sin k^{h}}{\sin k}+\frac{1}{2}\frac{\sin k}{\sin k^{h}}-\frac{\sin(k^{h}-k)}{k^{h}-k},\quad S_{2}:=-\frac{(\sin k^{h})\cos k}{2k}-\frac{(\sin k)\cos k^{h}}{2k^{h}}+\frac{\sin(k+k^{h})}{k+k^{h}},$ (12)
where $S_{1}$ mainly accounts for the dispersion error, and $S_{2}$ is small for large $k$. Note that
@@
Let $k^{h}:=\frac{2}{h}\arcsin\frac{kh}{2}$ and $c>0$. If $0<\frac{kh}{2}<1$, then $\frac{1}{24}k^{3}h^{2}<k^{h}-k<\frac{\pi-2}{8}k^{3}h^{2}$. If in addition $\sigma_{k}:=\min_{\xi\in\pi\mathbb{N}}|k-\xi|>0$ and $k^{3}h^{2}<\frac{8c\sigma_{k}}{\pi-2}$, then $k^{h}-k<c\sigma_{k}$ and $|\sin k^{h}|>\frac{2(1-c)}{\pi}\sigma_{k}$.
@@
Since $x+\frac{x^{3}}{6}<\arcsin x<x+(\frac{\pi}{2}-1)x^{3}$ for $0<x<1$, we have $k\!+\!\frac{1}{24}k^{3}h^{2}<k^{h}<k\!+\!\frac{\pi\!-\!2}{8}k^{3}h^{2}<k\!+\!c\sigma_{k}$. It follows that $\min_{\xi\in\pi\mathbb{N}}|\xi\!-\!k^{h}|>\min_{\xi\in\pi\mathbb{N}}|\xi\!-\!k|-|k^{h}\!-\!k|>(1\!-\!c)\sigma_{k}$. So $|\sin k^{h}|\geq\frac{2(1\!-\!c)}{\pi}\sigma_{k}$. ∎
@@
Under the assumptions of Lemma 2 with $0<c<1$, for $S_{1}$ in (13) it holds that
@@
Let $k\in((n_{k}^{+}-1)\pi,n_{k}^{+}\pi)$ for an integer $n_{k}^{+}$. By Lemma 2, it holds that $0<k^{h}-k<c\sigma_{k}<\frac{\pi}{2}$. Moreover, $(n_{k}^{+}-1)\pi+\sigma_{k}\leq k<k^{h}<k+c\sigma_{k}\leq n_{k}^{+}\pi-(1-c)\sigma_{k}$. Hence $(n_{k}^{+}-1)\pi<k<k^{h}<n_{k}^{+}\pi$, and $\cot k-\cot k^{h}=\frac{k^{h}-k}{\sin^{2}\tilde{k}}>k^{h}-k$ for some $\tilde{k}\in(k,k^{h})$. Since
[…]
@@
$S_{1}<\frac{1}{2}\frac{(k^{h}-k)^{2}}{\sin^{2}k}+1-\left(1-\frac{1}{6}(k^{h}-k)^{2}\right).$
@@
Let $k^{h}:=\frac{2}{h}\arcsin\frac{kh}{2}$ and $S_{2}$ given in (14). If $0<\frac{kh}{2}<1$, then it holds that
$\left(\frac{1}{4k^{2}}+\frac{1}{4k^{3}}\right)(k^{h}-k)^{2}<S_{2}<\left(\frac{1}{4k^{2}}+\frac{1}{4k^{3}}\right)(k^{h}-k)^{2}.$
@@
$|u\|\geq\left||g_{1}|\cdot\left\|\frac{\sin(kx)}{\sin k}\right\|-|g_{0}|\cdot\left\|\frac{\sin(k(1-x))}{\sin k}\right\|\right$|
@@
Under the assumptions of Lemma 2 with $0<c<1$, for $\tilde{S}_{1}$ in (17) it holds that
@@
Let $k^{h}:=\frac{2}{h}\arcsin\frac{kh}{2}$ and $\tilde{S}_{2}$ be given in (18). If $0<\frac{kh}{2}<1$, then it holds that
$\left(\frac{1}{4k^{2}}+\frac{1}{4k^{3}}\right)(k^{h}-k)^{2}<\tilde{S}_{2}<\left(\frac{1}{4k^{2}}+\frac{1}{4k^{3}}\right)(k^{h}-k)^{2}.$